OpenStudy (meilendurcer):

A bag contains 2 white marbles and 7 purple marbles. Two marbles are drawn at random. One marble is drawn and not replaced. Then a second marble is drawn. What is the probability of selecting a purple marble and then a white marble? What is the probability of selecting two white marbles? Is there a greater chance of selecting two white marbles in a row or two purple marbles in a row? Show your work.

OpenStudy (anonymous):

first is purple, probability is \(\frac{7}{9}\) second is white given the first is purple probability is \(\frac{2}{8}\) both happen, multiply the probabilities \[\frac{7}{9}\times \frac{2}{8}\]

OpenStudy (anonymous):

@satellite73 hi!! i've been waiting for somebody to help me all night, can you pleaseee help me when you're done here? :)

OpenStudy (anonymous):

two white marbles: probability first is white is \(\frac{2}{9}\) probability second is white given that the first is white is \(\frac{1}{8}\) again probability both happen is \[\frac{2}{9}\times \frac{1}{8}\]

OpenStudy (anonymous):

third one is rather obvious, since there are many more purple marbles than white ones

OpenStudy (anonymous):

@erin512 repost and i will look

OpenStudy (meilendurcer):

I am done thanks.

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

@satellite73 just reposted